# Evolving Language Models without Labels: Majority Drives Selection, Novelty Promotes Variation

### v2

@@ Abstract
Large language models (LLMs) are increasingly trained with reinforcement learning from verifiable rewards (RLVR), yet real-world deployment demands models that can self-improve without labels or external judges. Existing self-improvement approaches primarily rely on self-confirmation signals (e.g., confidence, entropy, or consistency) to generate rewards. This reliance drives models toward over-confident, majority-favored solutions, causing an entropy collapse that degrades pass@$n$ and reasoning complexity. To address this, we propose Evol-RL, a label-free framework that mirrors the evolutionary principle of balancing selection with variation. Concretely, Evol-RL retains the majority-voted answer as an anchor for stability, but adds a novelty-aware reward that scores each sampled solution by how different its reasoning is from other concurrently generated responses. This majority-for-stability + novelty-for-exploration rule mirrors the variation–selection principle: selection prevents drift, while novelty prevents collapse. Evaluation results show that Evol-RL consistently outperforms the majority-only baseline; e.g., training on label-free AIME24 lifts Qwen3-4B-Base AIME25 pass@1 from baseline’s 4.6% to 16.4%, and pass@16 from 18.5% to 37.9%. Evol-RL not only prevents in-domain diversity collapse but also improves out-of-domain generalization (from math reasoning to broader tasks, e.g., GPQA, MMLU-Pro, and BBEH).
@@ 1 Introduction
The reasoning capabilities of Large Language Models (LLMs) have advanced dramatically, particularly through paradigms like Reinforcement Learning with Verifiable Rewards (RLVR) (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025). The next frontier of intelligence lies in enabling LLMs to autonomously evolve, continuously learning from the vast, unlabeled data streams they encounter in real-world environments. This label-free evolving paradigm allows a model to iteratively improve itself while solving tasks, without relying on ground-truth labels or external judges, making it both practical and necessary. However, turning inference into learning reopens a long-standing RL problem: balancing exploration and exploitation. This dilemma becomes especially severe in label-free settings, where models must rely on internal signals (e.g., inherent self-consistency, entropy, or confidence) to generate rewards for themselves (Grandvalet & Bengio, 2004; Lee et al., 2013; Zuo et al., 2025; Shafayat et al., 2025; Li et al., 2025b).
@@
The fundamental flaw in relying on internal signals is not merely that they are initially noisy or biased, but that the learning process itself actively degrades the quality of the reward signal over time (Liang et al., 2024; Zeng et al., 2024). By rewarding conformity to its self-confirmation, the model systematically eliminates the solution diversity (Lee et al., 2024). This creates a degenerative feedback loop: a progressively narrower and more biased policy generates an increasingly impoverished reward signal, which in turn accelerates the policy’s collapse into a low-entropy state (Ding et al., 2025; Liang et al., 2024). Similar dynamics are well known in RL and self-training when entropy regularization or external supervision is absent (Haarnoja et al., 2018). Recent studies also show that training on self-generated data can harm diversity over time (Shumailov et al., 2024) and eventually lead to collapse. Figure 1 illustrates this phenomenon in reasoning: under traditional Test-Time Reinforcement Learning (TTRL) (Zuo et al., 2025), pass@1 may rise but pass@$n$ drops, while response length and complexity steadily decline, indicating that the model fails to evolve.
In this paper, we ground LLM evolving in the simple rule behind biological evolution: variation creates new candidates; selection keeps what works. Existing methods effectively implement only the selection half of evolution, driving the population toward whatever the model already believes. This majority-only (or entropy minimization, confidence maximization) reinforcement amplifies existing biases and often leads entropy collapse and shrinking response diversity, as shown above. Our formulation restores the full evolutionary loop: we pair selection, which stabilizes optimization by keeping high-quality solutions, with variation, which explicitly promotes novelty and sustains exploration. This idea is deeply rooted in decades of evolutionary computation research, including genetic algorithms (Holland, 1992; Eiben & Smith, 2015), novelty search (Lehman & Stanley, 2011), and quality–diversity (QD) methods such as MAP-Elites (Pugh et al., 2016), which collectively show that diversity preservation is essential for avoiding collapse and enabling robust, long-term progress.
@@
Contributions. (1) We diagnose why majority-only objectives shrink exploration during label-free training and formalize their link to entropy collapse on reasoning tasks. (2) We provide a new perspective on label-free learning by framing it as an evolutionary system. This view allows us to diagnose diversity collapse as a form of premature convergence and solve it by applying the core evolutionary principle of balancing selection with variation. (3) We design a practical novelty-aware reward that complements majority selection and enables stable, label-free improvement. Across math benchmarks, Evol-RL reverses the pass@$n$ decline, maintains longer and more informative chains of thought, and improves out-of-domain accuracy, while remaining simple to implement. (4) We deliver state-of-the-art results in unsupervised RL training, demonstrating that Evol-RL achieves significant out-of-domain generalization gains where prior methods fail, such as more than tripling pass@1 accuracy and doubling pass@16 accuracy on AIME25 benchmark.
2 Related Work
Enhancing Reasoning in Large Language Models.
Significant progress in LLM reasoning has been driven by RLVR (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025; Yu et al., 2025; Xiong et al., 2025; Dai et al., 2025b), which fine-tunes models using RL on tasks where an automated verifier can confirm the correctness of the final answer, such as mathematics and coding (Zeng et al., 2025; Wang et al., 2025a; b; Cui et al., 2025; Huang et al., 2025; Dai et al., 2025a; Zheng et al., 2025b; Zhou et al., 2025b; Zheng et al., 2025a; Fang et al., 2025). While highly effective, the reliance of RLVR on external verifiers restricts its applicability to domains with deterministic, easily checkable solutions (Zhao et al., 2025b; Zhou et al., 2025a; 2024). Our work contributes to the effort of improving reasoning in more general domains where such verifiers are unavailable.
@@ 2 Related Work
To overcome the limitations of verifiers and adapt to new data distributions, researchers have focused on label-free learning methods that generate reward signals without ground-truth labels. These approaches primarily fall into two categories. One line of research derives rewards from the model’s intrinsic confidence, training the model to become more "certain" by rewarding low-entropy or self-consistent outputs (Prabhudesai et al., 2025; Agarwal et al., 2025; Zhao et al., 2025a; Zhang et al., 2025b; Shafayat et al., 2025; Chung et al., 2025; Li et al., 2025a; Zhang et al., 2025a). The other prominent paradigm, which our work directly addresses, bootstraps supervision from majority. Test-Time Reinforcement Learning (TTRL) exemplifies this by using the majority-voted answer from multiple samples as a pseudo-label for RL updates (Zuo et al., 2025). While empirically powerful, we identify a critical flaw in the majority-driven approach: it suppresses solution diversity and actively punishes correct but non-mainstream reasoning, leading to the entropy collapse we describe. While ETTRL adjusts exploration within the original self-consistency framing (Liu et al., 2025), we are the first to pin down the majority trap and redesign the learning target to couple majority with population-level diversity.
@@ 3.1 Optimization with GRPO
For a given prompt $\mathbf{q}$, a policy LLM $\pi_{\theta_{\text{old}}}$ generates a group of $G$ complete responses $\{\mathbf{o}_{1},\ldots,\mathbf{o}_{G}\}$. Each response $\mathbf{o}_{i}$ receives a scalar reward $r_{i}$. Rewards within the group are normalized with a z-score to obtain a response-level advantage:
@@
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\!\left\{\frac{\pi_{\theta}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)}{\pi_{\theta_{\mathrm{old}}}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)}\,\hat{A}_{i,t},\;\operatorname{clip}\!\left(\frac{\pi_{\theta}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)}{\pi_{\theta_{\mathrm{old}}}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)},\,1-\epsilon_{\mathrm{low}},\,1+\epsilon_{\mathrm{high}}\right)\hat{A}_{i,t}\right\}$ (1)
@@ 3.2 Reward Design: Implementing Selection and Variation
[…]
@@
$\mathcal{L}_{\mathrm{ent}}(\theta)=-\lambda_{\mathrm{ent}}\;\mathbb{E}_{o\sim\pi_{\theta}}\left[\frac{1}{|o|}\sum_{t=1}^{|o|}H\!\left(\pi_{\theta}(\cdot\mid o_{<t},x)\right)\right],\qquad H(p)=-\sum_{v}p(v)\log p(v).$ (2)
@@
To prevent this, our three-part Variation strategy creates a robust exploratory dynamic. The entropy regularizer acts like a higher "mutation rate," ensuring a constant supply of diverse solutions for the system to work with. The novelty reward then provides directional pressure to this variation, giving a "survival bonus" to solutions that are semantically different from their peers. Finally, asymmetric clipping ensures that when a highly beneficial "mutation"—a rare, novel, and correct solution—appears, its strong learning signal is fully preserved for the next generation.
This design makes a collapsed state inherently unstable. In a uniform population of near-duplicate solutions, any "mutation" is, by definition, highly novel and receives a higher reward. The learning algorithm is thus forced to shift probability away from the uniform cluster and toward this more promising, distinct solution, ensuring that the policy remains robustly diverse.
@@ 4.1 Experimental Setup
Benchmarks. To test our method at scale, we use the large, standard MATH training set (MATH-TRAIN) (Hendrycks et al., 2021). We also follow the TTRL (Zuo et al., 2025) by training on two much smaller test sets: the general-purpose MATH-500 and the competition-level AIME24 (Li et al., 2024). This comprehensive setup allows us to validate Evol-RL’s versatility across both large-scale and specialized training conditions. Critically, during all training runs, we use only the problem statements, without any ground-truth labels or solutions. For evaluation, we assess the performance of our trained models on a diverse set of five benchmarks to measure both in-domain and out-of-domain generalization. The evaluation suite includes AIME24, AIME25, MATH500, AMC (Li et al., 2024), and GPQA-Diamond (GPQA) (Rein et al., 2024). Detailed training configuration can be found in Appendix A.
@@ 4.2 Main Results
The main results of our experiments are summarized in Table 1. We highlight four key findings that demonstrate the superiority of Evol-RL over the majority-only TTRL baseline.
Evol-RL Enhances Both Pass@1 and Pass@16 Performance. Across all experimental settings, Evol-RL consistently and substantially improves ‘pass@16‘ performance over TTRL, with gains frequently exceeding 20 percentage points on the most challenging benchmarks (e.g., +24.2pp on AIME24 for the 4B model). Evol-RL also delivers more consistent and substantial improvements to pass@1 accuracy than TTRL. This demonstrates that our method strengthens not only the model’s single-shot accuracy but also its ability to explore through multiple attempts.
@@
Generalization on Non-Mathematical Reasoning Tasks. The advantages of Evol-RL extend beyond the domain of mathematics. On the GPQA benchmark, where TTRL consistently causes pass@16 performance to degrade compared to the base model, Evol-RL reliably recovers and surpasses the base model. Across all training configurations, it achieves gains of +7 to +15 pp in pass@16 over TTRL, indicating that our diversity-preserving reward mechanism fosters a more generalizable reasoning ability that transfers effectively across different domains.
@@ 4.3 Ablation Study
Table 2: Performance of Qwen3-4B-Base with Evol-RL and its ablations on five benchmarks. Each cell reports pass@1/pass@16 accuracy.

### v3

@@ Abstract
Large language models (LLMs) are increasingly trained with reinforcement learning from verifiable rewards (RLVR), yet real-world deployment demands models that can self-improve without labels or external judges. Existing self-improvement approaches primarily rely on self-confirmation signals (e.g., confidence, entropy, or consistency) to generate rewards. This reliance drives models toward over-confident, majority-favored solutions, causing an entropy collapse that degrades pass@$n$ and reasoning complexity. To address this, we propose Evol-RL, a label-free framework that mirrors the evolutionary principle of balancing selection with variation. Concretely, Evol-RL retains the majority-voted answer as an anchor for stability, but adds a novelty-aware reward that scores each sampled solution by how different its reasoning is from other concurrently generated responses. This majority-for-stability + novelty-for-exploration rule mirrors the variation–selection principle: selection prevents drift, while novelty prevents collapse. Evaluation results show that Evol-RL consistently outperforms the majority-only baseline; e.g., training on label-free AIME24 lifts Qwen3-4B-Base AIME25 pass@1 from baseline’s 4.6% to 16.4%, and pass@16 from 18.5% to 37.9%. Evol-RL not only prevents in-domain diversity collapse but also improves out-of-domain generalization (from math reasoning to broader tasks, e.g., MMLU-Pro and BBEH).
@@ 1 Introduction
The reasoning capabilities of Large Language Models (LLMs) have advanced dramatically, particularly through paradigms like Reinforcement Learning with Verifiable Rewards (RLVR) (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025). The next frontier of intelligence lies in enabling LLMs to autonomously evolve, continuously learning from the vast, unlabeled data streams they encounter in real-world environments. This label-free evolving paradigm allows a model to iteratively improve itself while solving tasks, without relying on ground-truth labels or external judges, making it both practical and necessary. However, turning inference into learning reopens a long-standing RL problem: balancing exploration and exploitation. This dilemma becomes especially severe in label-free settings, where models must rely on internal signals (e.g., inherent self-consistency, entropy, or confidence) to generate rewards for themselves (Grandvalet and Bengio, 2004; Lee and others, 2013; Zuo et al., 2025; Shafayat et al., 2025; Li et al., 2025b).
@@
The fundamental flaw in relying on internal signals is not merely that they are initially noisy or biased, but that the learning process itself actively degrades the quality of the reward signal over time (Liang et al., 2024). By rewarding conformity to its self-confirmation, the model systematically eliminates the solution diversity (Lee et al., 2024). This creates a degenerative feedback loop: a progressively narrower and more biased policy generates an increasingly impoverished reward signal, which in turn accelerates the policy’s collapse into a low-entropy state (Ding et al., 2025; Liang et al., 2024; 2025b). Similar dynamics are well known in RL and self-training when entropy regularization or external supervision is absent (Haarnoja et al., 2018). Recent studies also show that training on self-generated data can harm diversity over time (Shumailov et al., 2024) and eventually lead to collapse. Figure 1 illustrates this phenomenon in reasoning: under traditional Test-Time Reinforcement Learning (TTRL) (Zuo et al., 2025), pass@1 may rise but pass@$n$ drops, while response length and complexity steadily decline, indicating that the model fails to evolve.
In this paper, we ground LLM evolving in the simple rule behind biological evolution: variation creates new candidates; selection keeps what works. Existing methods effectively implement only the selection half of evolution, driving the population toward whatever the model already believes. This majority-only (or entropy minimization, confidence maximization) reinforcement amplifies existing biases and often leads entropy collapse and shrinking response diversity, as shown above. Our formulation restores the full evolutionary loop: we pair selection, which stabilizes optimization by keeping high-quality solutions, with variation, which explicitly promotes novelty and sustains exploration. This idea is deeply rooted in decades of evolutionary computation research, including genetic algorithms (Holland, 1992; Eiben and Smith, 2015), novelty search (Lehman and Stanley, 2011), and quality–diversity (QD) methods such as MAP-Elites (Pugh et al., 2016), which collectively show that diversity preservation is essential for avoiding collapse and enabling robust, long-term progress.
@@
Contributions. (1) We diagnose why majority-only objectives shrink exploration during label-free training and formalize their link to entropy collapse on reasoning tasks. (2) We provide a new perspective on label-free learning by framing it as an evolutionary system. This view allows us to diagnose diversity collapse as a form of premature convergence and solve it by applying the core evolutionary principle of balancing selection with variation. (3) We design a practical novelty-aware reward that complements majority selection and enables stable, label-free improvement. Across math benchmarks, Evol-RL reverses the pass@$n$ decline, maintains longer and more informative chains of thought, and improves out-of-domain accuracy, while remaining simple to implement. (4) We deliver state-of-the-art results in unsupervised RL training, demonstrating that Evol-RL achieves significant out-of-domain generalization gains where prior methods fail, such as more than tripling pass@1 accuracy and doubling pass@16 accuracy on AIME25 benchmark. (5) We provide a theoretical analysis of entropy stabilization in Appendix D. We formally prove that while a correctness-only objective allows the optimal policy to collapse onto a single response, our novelty-augmented objective guarantees that the optimal policy must distribute probability mass across multiple correct modes, providing a rigorous foundation for the method’s stability.
2 Related Works
Enhancing Reasoning in LLMs.
Significant progress in LLM reasoning has been driven by RLVR (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025; Yu et al., 2025; Xiong et al., 2025; Dai et al., 2025b; Wang et al., 2025c; Zhuang et al., 2025), which fine-tunes models using RL on tasks where an automated verifier can confirm the correctness of the final answer, such as mathematics and coding (Zeng et al., 2025; Wang et al., 2025a; b; Cui et al., 2025; Huang et al., 2025; Dai et al., 2025a; Zheng et al., 2025b; Zhou et al., 2025b; Zheng et al., 2025a; Fang et al., 2025; Liu et al., 2025b). While highly effective, the reliance of RLVR on external verifiers restricts its applicability to domains with deterministic, easily checkable solutions (Zhao et al., 2025c; a; Zhou et al., 2025a; 2024; Liang et al., 2025a). Our work contributes to the effort of improving reasoning in more general domains where such verifiers are unavailable.
@@ 2 Related Works
To overcome the limitations of verifiers and adapt to new data distributions, researchers have focused on label-free learning methods. These approaches primarily fall into two categories. One line of research derives rewards from the model’s intrinsic confidence, training the model to become more "certain" by rewarding low-entropy or self-consistent outputs (Prabhudesai et al., 2025; Agarwal et al., 2025; Zhao et al., 2025b; Zhang et al., 2025b; Shafayat et al., 2025; Chung et al., 2025; Zhang et al., 2025a). The other prominent paradigm, which we directly address, bootstraps supervision from majority, exemplified by TTRL (Zuo et al., 2025). While empirically powerful, we identify a critical flaw in the majority-driven approach: it suppresses solution diversity and actively punishes correct but non-mainstream reasoning, leading to entropy collapse. Crucially, generic strategies like adding entropy loss or clip-high (Cui et al., 2025; Park et al., 2025) are insufficient to escape this “majority trap.” Instead, we propose a directional novelty reward that re-ranks credit based on semantic uniqueness. This fully label-free approach fundamentally redesigns the reward signal, distinguishing our work from methods that separate trained evaluators (Li et al., 2025a; Pang et al., 2023) or simple exploration adjustments (Liu et al., 2025a).
@@ 3.1 Optimization with GRPO
For a prompt $\mathbf{q}$, a policy LLM $\pi_{\theta_{\text{old}}}$ generates a group of $G$ responses $\{\mathbf{o}_{1},\ldots,\mathbf{o}_{G}\}$. Each response $\mathbf{o}_{i}$ receives a scalar reward $r_{i}$. Rewards within the group are normalized with a z-score to obtain a response-level advantage:
@@
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\!\Bigg\{\frac{\pi_{\theta}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)}{\pi_{\theta_{\mathrm{old}}}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)}\,\hat{A}_{i,t},$ (1)
$\operatorname{clip}\!\left(\frac{\pi_{\theta}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)}{\pi_{\theta_{\mathrm{old}}}\!\big(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t}\big)},\,1-\epsilon_{\mathrm{low}},\,1+\epsilon_{\mathrm{high}}\right)\hat{A}_{i,t}\Bigg\}$
@@ 3.2 Reward Design: Implementing Selection and Variation
$r_{i}\;=\;\begin{cases}-1,&\text{if invalid};\\[2.0pt] 0.5+0.5\,\tilde{u}_{i}\in[0.5,1],&\text{if }y_{i}=+1\ ;\\[2.0pt] -1+0.5\,\tilde{u}_{i}\in[-1,-0.5],&\text{if }y_{i}=-1\ .\end{cases}$
@@
$\mathcal{L}_{\mathrm{ent}}(\theta)$ $=-\lambda_{\mathrm{ent}}\;\mathbb{E}_{o\sim\pi_{\theta}}\left[\frac{1}{|o|}\sum_{t=1}^{|o|}H\!\left(\pi_{\theta}(\cdot\mid o_{<t},x)\right)\right],$ (2)
$H(p)$ $=-\sum_{v}p(v)\log p(v).$
@@
To prevent this, our three-part Variation strategy creates a robust exploratory dynamic. The entropy regularizer acts like a higher "mutation rate," ensuring a constant supply of diverse solutions for the system to work with. The novelty reward then provides directional pressure to this variation, giving a “survival bonus” to solutions that are semantically distinct from their peers. Finally, asymmetric clipping ensures that when a highly beneficial "mutation"—a rare, novel, and correct solution—appears, its strong learning signal is fully preserved for the next generation.
This design makes a collapsed state inherently unstable. In a uniform population, any novel solution receives a higher reward, forcing the algorithm to shift probability toward diverse solutions.
Theoretical Support. We formally validate this evolutionary intuition in Appendix D. We prove that while a correctness-only objective ($J_{0}$) admits degenerate optimal policies that collapse onto a single solution, our diversity-augmented objective ($J_{\lambda}$) alters the optimization landscape. Under mild assumptions, the global maximizer of our objective is theoretically guaranteed to spread probability mass across all correct reasoning modes (Theorem 1), thereby preventing the entropy collapse observed in standard self-training.
@@ 4.1 Experimental Setup
Benchmarks. To test our method at scale, we use the large, standard MATH training set (MATH-TRAIN) (Hendrycks et al., 2021). We also follow the TTRL (Zuo et al., 2025) by training on two much smaller test sets: the general-purpose MATH-500 and the competition-level AIME24 (Li et al., 2024). This comprehensive setup allows us to validate Evol-RL’s versatility across both large-scale and specialized training conditions. Critically, during all training runs, we use only the problem statements, without any ground-truth labels or solutions. For evaluation, we assess the performance of our trained models on a diverse set of five benchmarks to measure both in-domain and out-of-domain generalization. The evaluation suite includes AIME24, AIME25, MATH500, AMC (Li et al., 2024), and GPQA-Diamond (GPQA) (Rein et al., 2024). Detailed training configuration can be found in Appendix A. Furthermore, we provide additional experimental results on OctoThinker-8B-Hybrid-Base in Appendix B.1. More baseline comparisons, including Entropy Minimization (Agarwal et al., 2025) and Self-Consistency (Wang et al., 2023; Huang et al., 2023), are presented in Appendix B.2. The analysis of the extra time incurred by embedding computation and similarity measurement has been moved to Appendix B.5. A case study in Appendix B.6 shows that embedding similarity reliably captures differences in reasoning paths.
@@ 4.2 Main Results
The main results of our experiments are summarized in Table 4.2. We highlight four key findings that demonstrate the superiority of Evol-RL over the majority-only TTRL baseline.
Evol-RL Enhances Both Pass@1 and Pass@16 Performance. Across all experimental settings, Evol-RL consistently and substantially improves ‘pass@16‘ performance over TTRL, with gains frequently exceeding 20 percentage points on the most challenging benchmarks (e.g., +24.2% on AIME24 for the 4B model). Evol-RL also delivers more consistent and substantial improvements to pass@1 accuracy than TTRL. This demonstrates that our method strengthens not only the model’s single-shot accuracy but also its ability to explore through multiple attempts.
@@
Generalization on Non-Mathematical Reasoning Tasks. The advantages of Evol-RL extend beyond the domain of mathematics. On the GPQA benchmark, where TTRL consistently causes pass@16 performance to degrade compared to the base model, Evol-RL reliably recovers and surpasses the base model. Across all training configurations, it achieves gains of +7 to +15 % in pass@16 over TTRL, indicating that our diversity-preserving reward mechanism fosters a more generalizable reasoning ability that transfers effectively across different domains.
@@ 4.3 Ablation Study
Table 2: Performance of Qwen3-4B-Base with Evol-RL and its ablations. Each cell reports pass@1/pass@16 accuracy.
@@ Appendix B Additional Experimental Results
Table 6: Comparison of models trained with TTRL and Evol-RL on MATH-500 on OctoThinker-8B-Hybrid-Base. Each cell shows pass@1/pass@16 (averaged over 32 rollouts). $\Delta$ uses red (+) for positive and blue for negative values.
Model MATH AIME24 AIME25 AMC GPQA
[…]
Base Model 33.8/79.8 1.5/13.5 1.3/13.9 16.2/56.9 26.3/85.7
w/TTRL 63.8/76.4 2.8/10.8 2.1/11.0 27.9/54.7 31.9/71.5
[…]
$\Delta$ -0.6/+9.9 +6.2/+19.5 +5.1/+11.4 +6.2/+10.9 +1.3/+14.2
B.1 Effectiveness on Different Model Architectures
To verify that our approach is not limited to a single model family, we conducted an additional experiment applying EVOL-RL and TTRL to the OctoThinker-8B-Hybrid-Base model (Wang et al., 2025d), a different architecture from the Qwen3 series. We used the same label-free training setup on the MATH-500 dataset.
The results, presented in Table B, strongly confirm our core thesis. The TTRL baseline, when applied to OctoThinker, exhibits the classic symptoms of entropy collapse: while it significantly improves in-domain pass@1 accuracy (e.g., on MATH, from 33.8% to 63.8%), it fails to improve multi-path accuracy. In fact, pass@16 performance degrades on AIME24 (from 13.5% to 10.8%) and GPQA (from 85.7% to 71.5%) compared to the base model.
In sharp contrast, EVOL-RL successfully prevents this collapse and translates exploration into robust performance gains. While achieving a comparable pass@1 improvement on MATH, EVOL-RL yields massive improvements in pass@16 across all benchmarks. Most notably, it achieves a +19.5% gain on AIME24 and a +11.4% gain on AIME25 in pass@16 accuracy over TTRL.
This experiment demonstrates that entropy collapse is a fundamental flaw of the majority-only objective and that EVOL-RL is a robust and generalizable solution that functions effectively across different model architectures.
Table 7: Extended comparison of baseline methods on Qwen3-4B-Base, trained in a label-free setting using MATH-Train. Each cell shows pass@1/pass@16, and the highest value in each column is bolded.
Model MATH AIME24 AIME25 AMC GPQA
[…]
Base Model 67.4/89.6 10.0/32.4 5.5/30.0 39.3/75.2 34.4/85.6
w/TTRL 75.4/86.9 12.1/23.2 6.8/28.6 42.5/75.2 36.5/81.4
w/EM-RL-Token 76.0/90.6 12.5/31.3 10.5/30.8 46.6/77.7 36.8/82.6
w/EM-RL-Sequence 67.4/89.9 10.6/31.4 7.1/28.8 39.6/73.7 34.5/86.2
w/Self-Consistency 76.0/89.7 12.5/30.4 10.4/33.6 48.1/78.1 35.9/81.2
[…]
B.2 Comparison with Other Self-Improvement Baselines
To demonstrate that EVOL-RL is not just an improvement over TTRL but a more robust solution to the "entropy collapse" problem, we compare it against a broader suite of label-free self-improvement methods. These include methods based on self-consistency (Self-Consistency (Wang et al., 2023; Huang et al., 2023)) and intrinsic confidence (EM-RL-Token and EM-RL-Sequence (Agarwal et al., 2025)). We trained all methods on the Qwen3-4B-Base model under the same label-free setting, with results presented in Table B.1.
The results highlight a clear and consistent pattern: methods that optimize for a single signal (like consensus or confidence) fail to achieve robust, generalizable gains.
While baselines like TTRL, EM-RL-Token, and Self-Consistency all show moderate improvements in pass@1 accuracy on some benchmarks, they don’t show any consistent improvement in pass@16 performance, which is an indicator of entropy collapse. On the challenging AIME24 benchmark, every single one of these baselines performs worse than the original Base Model on pass@16 (e.g., 23.2% for TTRL and 30.4% for Self-Consistency, vs. 32.4% for the Base Model). This strongly suggests that their singular focus on "certainty" actively degrades solution diversity and multi-path reasoning.
In stark contrast, Evol-RL is the only method that robustly improves both single-shot accuracy (pass@1) and multi-path reliability (pass@16) across all benchmarks. The gains are most pronounced on out-of-domain tasks. On AIME24, Evol-RL achieves a pass@1 of 20.7% (vs. $\sim$12.5% for the next-best baselines) and a pass@16 of 47.6%, demonstrating a massive +15% improvement over even the Base Model, whereas all other methods failed. This result strongly supports our central thesis: a simple consensus or confidence signal is insufficient. True self-improvement requires an explicit mechanism—like our majority-for-stability + novelty-for-exploration rule—to prevent entropy collapse and foster diverse, generalizable reasoning.